# Positive mass and isoperimetry for continuous metrics with nonnegative scalar curvature

### v2

@@ 4.1. Producing a set satisfying the reverse Euclidean isoperimetric inequality
appearing in Theorem 3.1 (and defined in (3.2), (3.9), (3.10)), are uniformly bounded from above by a universal constant multiplied by the value of the corresponding constants on $\mathbb{R}^{3}$, which are independent of $\rho$. Moreover, since $F_{x}:(B^{i}_{16\rho+16}(x),g_{i})\to(\mathbb{R}^{3},g_{\mathrm{eu}})$ is $(1+3\eta)$-biLipschitz with its image, and contains the ball $B_{8\rho+8}^{\mathbb{R}^{3}}(0)$, the following holds: for every $r\leq\rho+1$ we can connect any two points $p,q\in\partial B^{i}_{r}(o)$ with a continuous curve $\gamma$ in the annulus $A^{i}_{3r/4,5r/4}(o)$. Indeed, it suffices to connect $F_{x}(p)$ with $F_{x}(q)$ with a curve $\tilde{\gamma}\subset\mathbb{R}^{3}$ in the annulus $A^{\mathbb{R}^{3}}_{7r/8,9r/8}(0)$, and take $\gamma:=F_{x}^{-1}(\tilde{\gamma})$. Moreover, with the same reasoning, the constant $C_{\mathrm{cov}}(B_{\rho+1}(o))$ defined in (3.11) is bounded above by a universal constant which only depends on the following universal constant independent of $\rho$:
$\begin{split}\tilde{C}\vcentcolon=\min\Big{\{}N\in\mathbb{N}\ :\ &\text{$A^{\mathbb{R}^{3}}_{r/2,3r/2}(0)$ is covered by $N$ open balls of radius $r/4$}\\ &\text{with centers in $A^{\mathbb{R}^{3}}_{r/2,3r/2}(o)$ for any $0<r\leq\rho+1$}\Big{\}}.\end{split}$ (4.9)
Hence, we can apply Theorem 3.1 on the ball $B^{i}_{2R}(o):=B^{i}_{4\rho+4}(o)$, where the radius $\rho+1$ in this proof corresponds to the number $\rho$ in the statement of Theorem 3.1. As a result of the discussion above, the constant $C$ appearing in Theorem 3.1 when applied to the ball $B^{i}_{4\rho+4}(o)$ is bounded above by a universal constant $\xi>1$ independent of $i$, $\rho$, for $i$ large enough.
Now let $w_{i}$ be the weak IMCF issuing from $o$ in $B^{i}_{4\rho+4}(o)$, given by Theorem 3.1. Denote $E_{t}^{i}:=\{w_{i}<t\}$. Let $T_{\rho,\xi}:=2\log(\rho-1)-\xi-1$, and take $\rho>1$ such that
@@
Notice that the choice of $\rho$ only depends on $\mathscr{P}$ and on the universal constant $\xi$. By the last assertion of Theorem 3.1, applied with $r:=\rho-1$, we have that
@@
Now we aim at applying item (3) of Section 3.2 with the choices $U:=B_{4\rho+4}^{i}(o)$, $\Omega:=F_{o}^{-1}(B_{2\rho}^{\mathbb{R}^{3}}(0))$, and $t_{1}:=T_{\rho,\xi}$. We stress that $\{w_{i}<T_{\rho,\xi}\}\subset\subset F_{o}^{-1}(B_{2\rho}^{\mathbb{R}^{3}}(0))$, and
@@
Moreover, for any arbitrary $t_{0}<T_{\rho,\xi}$, one can define $\Omega^{\prime\prime}$ to be a sufficiently small ball so that all the hypotheses of item (2) in Section 3.2 are met, since we have $w_{i}(x)\to-\infty$ as $x\to o$. Finally noticing that $H_{1}(F_{o}^{-1}(B_{2\rho}^{\mathbb{R}^{3}}(0));\mathbb{Z})=\{0\}$, one gets that all the hypotheses of item (2) and (3) in Section 3.2 are met, and thus $\partial\{w_{i}<t\}$ is connected for every $t<T_{\rho,\xi}$.
Therefore, recalling that $R_{g_{i}}\geq-\varepsilon_{i}$ on $B^{i}_{4\rho+4}(o)$, by Section 4.1 we get
@@
where we used that $P_{i}(E_{t}^{i})=4\pi e^{t}$, see Section 3.1, where $P_{i}(\cdot)$ denotes perimeter computed with respect to the metric $g_{i}$.
Now, since $\Omega^{i}_{T_{\rho,\xi}}\subset\subset B_{\rho}(o)$, and the perimeters $P_{i}(\Omega^{i}_{T_{\rho,\xi}})$ are equibounded, we can use the precompactness and lower semicontinuity result in Section 2.3 to get a set $E_{\rho}\subset B_{\rho}(o)$ such that $\Omega^{i}_{T_{\rho,\xi}}\to E_{\rho}\subset\subset\Omega\setminus\mathcal{C}$ in $L^{1}$, $P(E_{\rho})\leq\mathscr{P}$, and, by passing (4.11) to the limit, such that
@@
Finally, notice that $E_{\rho}\subset\subset B_{\rho+1}(o)$. Moreover (4.7) holds. This implies that, if $\eta<1$ is small enough, on $B_{\rho+1}(o)$ there holds a $(1,1^{*})$-Sobolev inequality (3.2) with $C_{1,{\rm Sob}}(B_{\rho+1}(o))$ bounded from above by a universal constant only depending on the constant in the Euclidean $(1,1^{*})$-Sobolev inequality. By Section 3, this implies that $P(E)\geq\tilde{\vartheta}|E|^{2/3}$ for a universal $\tilde{\vartheta}$, for every $E\subset\subset B_{\rho+1}(o)$. Applying the latter inequality on $E_{\rho}$, and using again (4.12), we finally get $P(E_{\rho})\geq\vartheta\mathscr{P}^{2/3}$ for a universal $\vartheta$, concluding the proof. ∎
@@ 4.2. Proof of the main results and consequences
Proof of Theorem 1.1.
It is a direct consequence of Section 4.1 and the definition of quasi-local isoperimetric mass in Section 1.1. ∎
Proof of Theorem 1.2.
The proof follows from a simplification of the arguments that led to Section 4.1. Indeed, we can apply Section 4.1 to $B^{i}_{r(o)}(o)$, where $r(o)$ is chosen so that $B^{i}_{r(o)}(o)$ is contractible for any $i$ sufficiently large. As in the proof of Section 4.1, one can find a common time $T\in\mathbb{R}$ such that weak inverse mean curvature flows $w^{i}_{r(o)}$ in $B^{i}_{r(o)}(o)$ issuing from $o$ are well-defined with $\{w^{i}_{r(o)}\leq T\}\Subset B^{i}_{r(o)}(o)$. Passing to the limit as $i\to+\infty$ suitable sublevel sets $\{w^{i}_{r(o)}\leq t\}$, for $t<T$, one gets as in Section 4.1 limit sets $E_{t}$ satisfying the reverse isoperimetric inequality with sharp constant. By (3.13), one can also choose $t$ so that the sets $E_{t}$ are contained in any arbitrarily small ball centered at $o$. ∎
Proof of Theorem 1.3.
@@
Let $v\in(2v_{0},+\infty)$. By Theorem 2.2 there is an isoperimetric set $E\subset M$ with $|E|\leq v_{0}$ (possibly empty) and a ball $B\subset\mathbb{R}^{3}$ with $|B|\geq v-v_{0}\geq v/2$ such that
@@
Let now $\varepsilon>0$ be such $B_{\varepsilon}$ is a ball in $\mathbb{R}^{3}$ concentric to $B$ and with volume $|B_{\varepsilon}|_{\mathrm{eu}}=|B|_{\mathrm{eu}}-\varepsilon$. Notice that, by approximating $B_{\varepsilon}\subset\mathbb{R}^{3}$ with sets diverging along the manifold, arguing as in the proof of the upper bound for the isoperimetric profile in Section 2.3, we get $I(v-\varepsilon)\leq P(E)+P_{\mathrm{eu}}(B_{\varepsilon})$. Thus, taking $\varepsilon\to 0$, we find
@@
Since $I$ is continuous by Section 2.3, the latter implies that $I$ is strictly increasing on $(2v_{0},+\infty)$. Thus the sought claim is proved.
We aim now at showing that there exists an isoperimetric set with volume strictly greater than $2v_{0}$, thus reaching a contradiction. Fix $v>2v_{0}$. By Theorem 2.2 there is an isoperimetric set $E\subset M$ with $|E|\leq v_{0}$ and a ball $B\subset\mathbb{R}^{3}$ with $|B|\geq v-v_{0}>0$ such that
@@
By Theorem A.1 we know that $E$ is bounded.
We apply Section 4.1 with $K=\emptyset$ and $K^{\prime}=\overline{B}_{r}(o)$ for some ball $B_{r}(o)$ with $r>1$ such that $E\cup\mathcal{C}\subset B_{r-1}(o)$. Then there is a set $F\subset M$ such that $F\subset\subset M\setminus K^{\prime}$, $P(F)\leq P_{\mathrm{eu}}(B)$, and
@@
Hence all the inequalities in the previous formula are equalities, and then $E\cup F$ is an isoperimetric set with volume $>2v_{0}$, which is a contradiction.
We now briefly argue for the existence of isoperimetric sets with small volume. Assume again for a contradiction that there exists $v_{0}$ such that no isoperimetric set of volume $v$ exists for any $v\in(0,v_{0})$. Then, by Theorem 2.2, for every $v\in(0,v_{0})$ we have $I(v)=P_{\rm eu}(B_{v})$, where $B_{v}$ is the round ball in $(\mathbb{R}^{3},\delta)$ of volume $v$. We can now reach a contradiction as follows. By Theorem 1.2 there is $v^{\prime}\in(0,v_{0})$ and $F$ such that
@@
In the proof of Theorem 1.3 we argued that, if for some $v_{0}>0$ no isoperimetric sets exist for any volume $v>v_{0}$ then the isoperimetric profile is strictly increasing for large volumes. This implied that some isoperimetric set of large volume exists, resulting in a contradiction. In fact, if the isoperimetric profile is strictly increasing, then one can deduce that isoperimetric sets exist for any volume. This is what happens in a smooth asymptotically flat $3$-manifold of nonnegative scalar curvature that is complete with no closed minimal surfaces or endowed with a horizon boundary. For this reason, the following result can be seen as a generalization of the existence result for any volume obtained in [CarlottoChodoshEichmair, Proposition K.1] (see also [BFSurvey, Theorem 3.6]). Moreover, when in addition the isoperimetric profile diverges at infinity, such isoperimetric sets realize the isoperimetric mass in the sense of (4.13). The isoperimetric profile diverges at infinity for instance when a global Euclidean-like isoperimetric inequality is in force. The latter happens, for example, when the manifold is $C^{0}$-asymptotically flat globally.
[…]
@@
We assume by contradiction that, for some $v>0$, there exists no isoperimetric sets of volume $v$. By Theorem 2.2, we have $I(v)=P(E)+P_{\rm eu}(B)$ for a possibly empty $E\subset M$ realizing $P(E)=I(|E|)$ and $B\subset\mathbb{R}^{3}$ a nonempty ball, such that $|E|+|B|_{\rm eu}=v$. The proof of Theorem 1.3 shows that we can find a set $F\subset\subset M\setminus\overline{B}_{r+1}(o)$, for some ball such that $E\subset B_{r}(o)$, with $P(F)\leq P_{\rm eu}(B)$ and with volume $|F|\geq|B|_{\rm eu}$. If $|F|=|B|_{\rm eu}$, then $E\cup F$ is an isoperimetric set of volume $v$, giving a contradiction. If $\left\lvert F\right\rvert>\left\lvert B\right\rvert_{\rm eu}$, we derive a contradiction with strict monotonicity of $I$. Indeed, on the one hand we would have
@@
We close this section with the proof of Theorem 1.4.
Proof of Theorem 1.4.
Let $o\in M$ and $B_{R}(o)\Subset\Omega$ with $H_{1}(B_{R}(o),\mathbb{Z})=\{0\}$. Let $w$ be given by Theorem 3.1. By (1.5), any set compactly contained in $\Omega$ satisfies the Euclidean isoperimetric inequality. Hence Section 4.1 implies that (4.1) holds with $\delta=0$ and with the equality sign for all $t\in(-\infty,T)$ for some $T$, i.e., sublevel sets of the local weak IMCF from $o$ achieve equality in the Euclidean isoperimetric inequality.
Going back to the proof of Section 4.1, it follows that all the inequalities in (4.4) are equalities for all $t\in(-\infty,T)$. This implies that the Hawking mass $\mathfrak{m}_{H}(\partial E_{t})$ vanishes for all $t\in(-\infty,T)$. At this point, we can conclude following the argument in [HuiskenIlmanen, pp. 422-424]. For the reader’s convenience, we briefly include the line of reasoning.
Using that the derivative of the Hawking mass [HuiskenIlmanen, Geroch Monotonicity Formula 5.8] vanishes (see also [BenattiPludaPozzetta, Theorem 5.5] for the version obtained through $p$-harmonic approximation, as in the present case), one gets that $\partial E_{t}$ is a smooth totally umbilical hypersurface with constant mean curvature for any $t\in(-\infty,T)$. Also, the hypersurfaces $\partial E_{t}$ are not minimal since $\int_{\partial E_{t}}H^{2}=16\pi$ as $\mathfrak{m}_{H}(\partial E_{t})=0$. One can also deduce from the latter that the flow does not jump.
It follows that the hypersurfaces $\partial E_{t}$ yield a smooth foliation of a neighborhood of $o$. Denoting by $H(t)>0$ the constant mean curvature of $\partial E_{t}$, we can rewrite the metric $g$ on $\{w<T\}$ as $g|_{\{w<T\}}=H^{-2}(t)\mathrm{d}t\otimes\mathrm{d}t+g_{\partial E_{t}}$. Differentiating $g_{\partial E_{t}}$ with respect to $t$ and exploiting the total umbilicity of $\partial E_{t}$, the explicit value of $H^{2}(t)$, and the Euclidean asymptotic behavior of $g$ close to $o$, one obtains that ($(\{w<T\},g)$ is isometric to $\left([0,S)\times\mathbb{S}^{2},\mathrm{d}s\otimes\mathrm{d}s+r^{2}g_{\mathbb{S}^{2}}\right)$, for some $S>0$, where $g_{\mathbb{S}^{2}}$ is the standard round metric on the sphere. Thus, ($(\{w<T\},g)$ is flat, as desired. We refer the reader also to [benatti2022minkowski, Proof of Theorem 1.2] for additional details on this computation. ∎

### v3

@@ 4.1. Producing a set satisfying the reverse Euclidean isoperimetric inequality
[…]
Hence, we can apply Theorem˜3.6 on the ball $B^{i}_{2R}(o):=B^{i}_{4\rho+4}(o)$, where the radius $\rho+1$ in this proof corresponds to the number $\rho$ in the statement of Theorem˜3.6. As a result of the discussion above, the constant $C$ appearing in Theorem˜3.6 when applied to the ball $B^{i}_{4\rho+4}(o)$ is bounded above by a universal constant $\xi>1$ independent of $i$, $\rho$, for $i$ large enough.
Now let $w_{i}$ be the weak IMCF issuing from $o$ in $B^{i}_{4\rho+4}(o)$, given by Theorem˜3.6. Denote $E_{t}^{i}:=\{w_{i}<t\}$. Let $T_{\rho,\xi}:=2\log(\rho-1)-\xi-1$, and take $\rho>1$ such that
@@
Notice that the choice of $\rho$ only depends on $\mathscr{P}$ and on the universal constant $\xi$. By the last assertion of Theorem˜3.6, applied with $r:=\rho-1$, we have that
@@
Now we aim at applying item (3) of Section˜3.2 with the choices $U:=B_{4\rho+4}^{i}(o)$, $\Omega:=F_{o}^{-1}(B_{2\rho}^{\mathbb{R}^{3}}(0))$, and $t_{1}:=T_{\rho,\xi}$. We stress that $\{w_{i}<T_{\rho,\xi}\}\subset\subset F_{o}^{-1}(B_{2\rho}^{\mathbb{R}^{3}}(0))$, and
@@
Moreover, for any arbitrary $t_{0}<T_{\rho,\xi}$, one can define $\Omega^{\prime\prime}$ to be a sufficiently small ball so that all the hypotheses of item (2) in Section˜3.2 are met, since we have $w_{i}(x)\to$-$\infty$ as $x\to o$. Finally noticing that $H_{1}(F_{o}^{-1}(B_{2\rho}^{\mathbb{R}^{3}}(0));\mathbb{Z})=\{0\}$, one gets that all the hypotheses of item (2) and (3) in Section˜3.2 are met, and thus $\partial\{w_{i}<t\}$ is connected for every $t<T_{\rho,\xi}$.
Therefore, recalling that $R_{g_{i}}\geq-\varepsilon_{i}$ on $B^{i}_{4\rho+4}(o)$, by Section˜4.1 we get
@@
where we used that $P_{i}(E_{t}^{i})=4\pi e^{t}$, see Section˜3.1, where $P_{i}(\cdot)$ denotes perimeter computed with respect to the metric $g_{i}$.
Now, since $\Omega^{i}_{T_{\rho,\xi}}\subset\subset B_{\rho}(o)$, and the perimeters $P_{i}(\Omega^{i}_{T_{\rho,\xi}})$ are equibounded, we can use the precompactness and lower semicontinuity result in Section˜2.3 to get a set $E_{\rho}\subset B_{\rho}(o)$ such that $\Omega^{i}_{T_{\rho,\xi}}\to E_{\rho}\subset\subset\Omega\setminus\mathcal{C}$ in $L^{1}$, $P(E_{\rho})\leq\mathscr{P}$, and, by passing (4.11) to the limit, such that
@@
Finally, notice that $E_{\rho}\subset\subset B_{\rho+1}(o)$. Moreover (4.7) holds. This implies that, if $\eta<1$ is small enough, on $B_{\rho+1}(o)$ there holds a $(1,1^{*})$-Sobolev inequality (3.2) with $C_{1,{\rm Sob}}(B_{\rho+1}(o))$ bounded from above by a universal constant only depending on the constant in the Euclidean $(1,1^{*})$-Sobolev inequality. By Section˜3, this implies that $P(E)\geq\tilde{\vartheta}|E|^{2/3}$ for a universal $\tilde{\vartheta}$, for every $E\subset\subset B_{\rho+1}(o)$. Applying the latter inequality on $E_{\rho}$, and using again (4.12), we finally get $P(E_{\rho})\geq\vartheta\mathscr{P}^{2/3}$ for a universal $\vartheta$, concluding the proof. ∎
@@ 4.2. Proof of the main results and consequences
Proof of Theorem˜1.4.
It is a direct consequence of Section˜4.1 and the definition of quasi-local isoperimetric mass in Section˜1.1. ∎
Proof of Theorem˜1.5.
The proof follows from a simplification of the arguments that led to Section˜4.1. Indeed, we can apply Section˜4.1 to $B^{i}_{r(o)}(o)$, where $r(o)$ is chosen so that $B^{i}_{r(o)}(o)$ is contractible for any $i$ sufficiently large. As in the proof of Section˜4.1, one can find a common time $T\in\mathbb{R}$ such that weak inverse mean curvature flows $w^{i}_{r(o)}$ in $B^{i}_{r(o)}(o)$ issuing from $o$ are well-defined with $\{w^{i}_{r(o)}\leq T\}\Subset B^{i}_{r(o)}(o)$. Passing to the limit as $i\to+\infty$ suitable sublevel sets $\{w^{i}_{r(o)}\leq t\}$, for $t<T$, one gets as in Section˜4.1 limit sets $E_{t}$ satisfying the reverse isoperimetric inequality with sharp constant. By (3.13), one can also choose $t$ so that the sets $E_{t}$ are contained in any arbitrarily small ball centered at $o$. ∎
Proof of Theorem˜1.6.
@@
Let $v\in(2v_{0},+\infty)$. By Theorem˜2.16 there is an isoperimetric set $E\subset M$ with $|E|\leq v_{0}$ (possibly empty) and a ball $B\subset\mathbb{R}^{3}$ with $|B|\geq v-v_{0}\geq v/2$ such that
@@
Let now $\varepsilon>0$ be such $B_{\varepsilon}$ is a ball in $\mathbb{R}^{3}$ concentric to $B$ and with volume $|B_{\varepsilon}|_{\mathrm{eu}}=|B|_{\mathrm{eu}}-\varepsilon$. Notice that, by approximating $B_{\varepsilon}\subset\mathbb{R}^{3}$ with sets diverging along the manifold, arguing as in the proof of the upper bound for the isoperimetric profile in Section˜2.3, we get $I(v-\varepsilon)\leq P(E)+P_{\mathrm{eu}}(B_{\varepsilon})$. Thus, taking $\varepsilon\to 0$, we find
@@
Since $I$ is continuous by Section˜2.3, the latter implies that $I$ is strictly increasing on $(2v_{0},+\infty)$. Thus the sought claim is proved.
We aim now at showing that there exists an isoperimetric set with volume strictly greater than $2v_{0}$, thus reaching a contradiction. Fix $v>2v_{0}$. By Theorem˜2.16 there is an isoperimetric set $E\subset M$ with $|E|\leq v_{0}$ and a ball $B\subset\mathbb{R}^{3}$ with $|B|\geq v-v_{0}>0$ such that
@@
By Theorem˜A.2 we know that $E$ is bounded.
We apply Section˜4.1 with $K=\emptyset$ and $K^{\prime}=\overline{B}_{r}(o)$ for some ball $B_{r}(o)$ with $r>1$ such that $E\cup\mathcal{C}\subset B_{r-1}(o)$, where $\mathcal{C}$ is as in the statement. Then there is a set $F\subset M$ such that $F\subset\subset M\setminus K^{\prime}$, $P(F)\leq P_{\mathrm{eu}}(B)$, and
@@
Hence all the inequalities in the previous formula are equalities, and then $E\cup F$ is an isoperimetric set with volume $2v_{0}$, which is a contradiction.
We now briefly argue for the existence of isoperimetric sets with small volume. Assume again for a contradiction that there exists $v_{0}$ such that no isoperimetric set of volume $v$ exists for any $v\in(0,v_{0})$. Then, by Theorem˜2.16, for every $v\in(0,v_{0})$ we have $I(v)=P_{\rm eu}(B_{v})$, where $B_{v}$ is the round ball in $(\mathbb{R}^{3},\delta)$ of volume $v$. We can now reach a contradiction as follows. By Theorem˜1.5 there is $v^{\prime}\in(0,v_{0})$ and $F$ such that
@@
In the proof of Theorem˜1.6 we argued that, if for some $v_{0}>0$ no isoperimetric sets exist for any volume $v>v_{0}$ then the isoperimetric profile is strictly increasing for large volumes. This implied that some isoperimetric set of large volume exists, resulting in a contradiction. In fact, if the isoperimetric profile is strictly increasing, then one can deduce that isoperimetric sets exist for any volume. This is what happens in a smooth asymptotically flat $3$-manifold of nonnegative scalar curvature that is complete with no closed minimal surfaces or endowed with a horizon boundary. For this reason, the following result can be seen as a generalization of the existence result for any volume obtained in [23, Proposition K.1] (see also [12, Theorem 3.6]). Moreover, when in addition the isoperimetric profile diverges at infinity, such isoperimetric sets realize the isoperimetric mass in the sense of (4.13). The isoperimetric profile diverges at infinity for instance when a global Euclidean-like isoperimetric inequality is in force. The latter happens, for example, when the manifold is $C^{0}$-asymptotically flat globally.
Proposition 4.3.
@@
We assume by contradiction that, for some $v>0$, there exists no isoperimetric sets of volume $v$. By Theorem˜2.16, we have $I(v)=P(E)+P_{\rm eu}(B)$ for a possibly empty $E\subset M$ realizing $P(E)=I(|E|)$ and $B\subset\mathbb{R}^{3}$ a nonempty ball, such that $|E|+|B|_{\rm eu}=v$. The proof of Theorem˜1.6 shows that we can find a set $F\subset\subset M\setminus\overline{B}_{r+1}(o)$, for some ball such that $E\subset B_{r}(o)$, with $P(F)\leq P_{\rm eu}(B)$ and with volume $|F|\geq|B|_{\rm eu}$. If $|F|=|B|_{\rm eu}$, then $E\cup F$ is an isoperimetric set of volume $v$, giving a contradiction. If $\left\lvert F\right\rvert>\left\lvert B\right\rvert_{\rm eu}$, we derive a contradiction with strict monotonicity of $I$. Indeed, on the one hand we would have
@@
We close this section with the proof of Theorem˜1.7.
Proof of Theorem˜1.7.
Let $o\in M$ and $B_{R}(o)\Subset\Omega$ with $H_{1}(B_{R}(o),\mathbb{Z})=\{0\}$. Let $w$ be given by Theorem˜3.6. By (1.5), any set compactly contained in $\Omega$ satisfies the Euclidean isoperimetric inequality. Hence Section˜4.1 implies that (4.1) holds with $\delta=0$ and with the equality sign for all $t\in(-\infty,T)$ for some $T$, i.e., sublevel sets of the local weak IMCF from $o$ achieve equality in the Euclidean isoperimetric inequality.
Going back to the proof of Section˜4.1, it follows that all the inequalities in (4.4) are equalities for all $t\in(-\infty,T)$. This implies that the Hawking mass $\mathfrak{m}_{H}(\partial E_{t})$ vanishes for all $t\in(-\infty,T)$. At this point, we can conclude following the argument in [38, pp. 422-424]. For the reader’s convenience, we briefly include the line of reasoning.
Using that the derivative of the Hawking mass [38, Geroch Monotonicity Formula 5.8] vanishes (see also [14, Theorem 5.5] for the version obtained through $p$-harmonic approximation, as in the present case), one gets that $\partial E_{t}$ is a smooth totally umbilical hypersurface with constant mean curvature for any $t\in(-\infty,T)$. Also, the hypersurfaces $\partial E_{t}$ are not minimal since $\int_{\partial E_{t}}H^{2}=16\pi$ as $\mathfrak{m}_{H}(\partial E_{t})=0$. One can also deduce from the latter that the flow does not jump.
It follows that the hypersurfaces $\partial E_{t}$ yield a smooth foliation of a neighborhood of $o$. Denoting by $H(t)>0$ the constant mean curvature of $\partial E_{t}$, we can rewrite the metric $g$ on $\{w<T\}$ as $g|_{\{w<T\}}=H^{-2}(t)\mathrm{d}t\otimes\mathrm{d}t+g_{\partial E_{t}}$. Differentiating $g_{\partial E_{t}}$ with respect to $t$ and exploiting the total umbilicity of $\partial E_{t}$, the explicit value of $H^{2}(t)$, and the Euclidean asymptotic behavior of $g$ close to $o$, one obtains that $(\{w<T\},g)$ is isometric to $\left([0,S)\times\mathbb{S}^{2},\mathrm{d}s\otimes\mathrm{d}s+r^{2}g_{\mathbb{S}^{2}}\right)$, for some $S>0$, where $g_{\mathbb{S}^{2}}$ is the standard round metric on the sphere. Thus, $(\{w<T\},g)$ is flat, as desired. We refer the reader also to [10, Proof of Theorem 1.2] for additional details on this computation. ∎